OpenStudy (anonymous):

Find the partial derivative of z w.r.t. s: z=e^r * cos(k) where r = st and k = sqrt{s^2+t^2}

OpenStudy (anonymous):

Let's put the value of r and k \[z=e^{st}\times \cos (\sqrt {s^2+t^2})\] Treat every variable other than s as constant \[\frac{dz}{ds}=\cos (\sqrt {s^2+t^2})\frac d {ds} e^{st}+e^{st}\frac{d}{ds}\cos (\sqrt {s^2+t^2})\] Can you try @taljaards ?

OpenStudy (anonymous):

\[\frac{dz}{ds} = \cos(\sqrt{s^2+t^2}) \times s - e^{st}\sin(\sqrt{s^2+t^2}) \times 0.5(s^2+t^2)^{-0.5}2t\] ?

OpenStudy (anonymous):

derivative of st is t

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

Hi! Thanks for the help! I'm glad I studied hard for the test - a similar question to this was actually asked!